# Evaluating and Enhancing LLMs for Multi-turn Text-to-SQL with Multiple Question Types

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have significantly advanced text-to-SQL systems. However, most LLM-based methods often narrowly focus on SQL generation, neglecting the complexities of real-world conversational queries. This oversight can lead to unreliable responses, particularly for ambiguous questions that cannot be directly addressed with SQL. To bridge this gap, we propose MMSQL, a comprehensive test suite designed to evaluate the question classification and SQL generation capabilities of LLMs by simulating real-world scenarios with diverse question types and multi-turn Q&A interactions. Using MMSQL, we assessed the performance of popular LLMs, including both open-source and closed-source models, and identified key factors impacting their performance in such scenarios. Moreover, we introduce an LLM-based multi-agent framework that employs specialized agents to identify question types and determine appropriate answering strategies. Our experiments demonstrate that this approach significantly enhances the model’s ability to navigate the complexities of conversational dynamics, effectively handling the diverse and complex nature of user queries.
keywords:
@@
1 Introduction
Text-to-SQL bridges natural language and SQL queries, empowering non-technical users to access and analyze data without mastering complex SQL knowledge. The advent of LLMs, with their remarkable capacity for following instructions, has transformed the text-to-SQL domain. LLM-based methods have delivered remarkable outcomes across various text-to-SQL tasks [1, 2], while the assessment of their robustness is increasingly drawing attention. Current research often assumes that user queries are unambiguous, focusing on expanding the scope of databases and increasing the complexity of questions to elevate the difficulty of question-to-SQL mapping, thereby pushing the boundaries of text-to-SQL systems. [3, 4, 5, 6]. However, in practical scenarios, user queries are inherently dynamic and uncertain [7]. Assuming that all user queries can be answered may lead to the generation of hallucinated content and unreliable predictions, thereby undermining the dependability of these systems [8, 9]. Observing real-world conversations, it is common for users to engage in multi-turn dialogues with interrelated questions that do not readily translate into SQL [7, 10, 11]. Figure 1 illustrates examples of responses to three common types of such questions: (1) Unanswerable: If the database lacks data, such as the flight departure times from an airport queried in Turn 1, the system needs to explain its inability to provide an answer. (2) Ambiguous: In Turn 3, the query about the flight number of Delta Airlines might refer to conditions from Turn 1 and 2, ”from airport APG,” or it might simply mean all flight numbers operated by Delta Airlines. The system needs to recognize this ambiguity and respond accordingly to avoid providing incorrect SQL responses. (3) Improper: Questions unrelated to the database content, such as casual conversation in Turn 4, should not prompt SQL-based responses. These examples highlight the necessity for reliable models capable of navigating various types of complex conversational dynamics. There is a need for more robust models that can effectively manage such challenges.
@@ 1 Introduction
Existing studies have only sporadically addressed the effective assessment and handling of multi-type questions. Most text-to-SQL research focuses on achieving high accuracy for single-type or single-round user questions [4, 6, 12, 13], often overlooking the need to develop systems capable of multi-turn dialogues that handle a variety of question types [3]. This limited scope can result in incorrect responses, especially for those questions that do not necessitate SQL solutions. It has been observed that even the most sophisticated LLMs struggle to effectively address these challenges [9]. Although recent studies have acknowledged the challenges of ambiguous and unanswerable questions [7]. However, existing solutions often resort to simple abstention when faced with vague or conversational elements, lacking granular designs to manage various question types, leading to passive responses and hindering practical applicability and robustness [10, 14, 15]. In summary, while individual studies have made some progress in evaluating and enhancing the reliability of text-to-SQL systems, these efforts fall short of a comprehensive solution to the problem.
In response to these challenges, we introduce the Multi-type and Multi-turn text-to-SQL test suite (MMSQL), a comprehensive benchmark engineered to evaluate the proficiency of LLMs in handling multi-turn text-to-SQL tasks across diverse question types. Utilizing MMSQL, we have undertaken an exhaustive performance analysis of leading LLMs, encompassing open-source and closed-source models. This analysis provides detailed insights, highlighting the promise of open-source models and the significant benefits of fine-tuning. It also reveals the suboptimal performance of models on specific question types and the varying effectiveness of strategies designed to resolve ambiguity. We propose an innovative LLM-based multi-agent framework, inspired by top-performing models and informed by these insights. This framework is anchored by a core Question Detector and Question Decomposer, tasked with identifying question types and determining appropriate answering strategies. It could provide potential SQL queries for ambiguous questions, accounting for multiple interpretations. Furthermore, the framework includes two supportive agents: the Schema Selector, which identifies and provides the essential subset of a database schema, and the SQL Refiner, which is dedicated to refining SQL queries. Our experiments, conducted on the MMSQL benchmark, have shown that this framework yields a significant performance improvement.
@@
2 Related Work
2.1 Text-to-SQL
@@
Text-to-SQL involves transforming natural language queries into SQL, representing a crucial area of natural language processing with many practical applications. Crucial datasets include single-turn datasets like WikiSQL [16], Spider, and BIRD [3], as well as multi-turn datasets such as CHASE [13], SParC [12], and CoSQL [17]. These datasets serve as benchmarks for evaluating the performance of text-to-SQL systems. In recent years, the advent of pre-trained language models (PLMs) has driven significant progress in this domain. Models like SQLova [18], PICARD [19], and RAT-SQL [20] have achieved impressive results. However, despite their effectiveness, these specialized models require considerable resources and time for training. LLMs have revolutionized numerous NLP tasks, including text-to-SQL, surpassing previous models and establishing a new paradigm [21, 22]. Recent studies have focused on enhancing prompt design and developing complex multi-stage frameworks that leverage LLMs to boost performance [23]. Notably, LLM-based frameworks like DIN-SQL [24], which utilize chain-of-thought reasoning [25], decompose the problem into simpler sub-problems. MAC-SQL [6] and CHASE-SQL [5] employ a multi-agent collaborative framework, demonstrating significant advancements in this field.
@@
Recent works highlight natural language’s inherent diversity and ambiguity, noting that not all user queries can be effectively translated into accurate SQL as the answer in practical applications [7, 10]. Datasets such as TriageSQL [14], NoisySP [7], and CoSQL have incorporated question type detection as part of the model’s task. Previous work, TrustSQL [10] demands that models opt for abstention when faced with diverse unanswerable questions. Additionally, AMBROSIA [9] and [26] propose testing text-to-SQL systems by parsing ambiguous questions into multiple potential SQL queries, thereby enhancing model usability in real-world scenarios. [26] improves beam search on T5 [27] to provide possible SQL responses for ambiguous questions and evaluate their effectiveness. Additionally, [15] employs questions enhanced dialogue augmentation to train open-source LLMs to identify question types. These studies focus solely on the model’s ability to recognize question types or address specific problems, which may result in unhelpful abstentions or hallucinations due to specialized question-and-answer tailoring. Moreover, most work is limited to single-turn settings, whereas real interactions often involve ambiguity that must be addressed in multi-turn contexts. These limitations motivate our work, which evaluates and enhances LLMs for multi-turn dialogues with multiple question types.
2.2 LLM-based Agents
@@
LLM-based agents have long captured the attention of researchers in both academia and industry [28]. With the expansion of web knowledge, LLMs are increasingly demonstrating intelligence levels comparable to humans. This evolution has sparked a growing interest in creating autonomous agents driven by LLMs. AutoGPT [29] and MetaGPT [30] enhance AI models by integrating a range of useful tools, allowing developers to create adaptable, conversational agents that can function in various modes by leveraging LLMs, human input, and other tools to complete tasks. In the realm of text-to-SQL parsing, frameworks such as MAC-SQL [6], MAG-SQL [31], and CHASE-SQL [5] employ multiple LLM-based agents to collaboratively interpret SQL queries. This can typically be summarized as a three-step process: retrieving relevant values, generating SQL, and refining the SQL output. These multi-agent systems have achieved leading results on the BIRD benchmark [3]. Drawing inspiration from these advancements, we developed a multi-agent framework that incorporates a Question Detector Agent, designed to effectively manage the complexity and variety of user queries encountered in real-world applications.
3 MMSQL: Multi-type and Multi-turn Text-to-SQL Test Suite
3.1 Task Formulation
  1. 1.
    Ambiguous: These questions include terms that might correspond to multiple columns or cell values in the database, causing potential ambiguity in SQL query generation. The system should seek clarification from the user to resolve this ambiguity while simultaneously attempting to predict possible SQL queries. For example, if a term like ”Glenn” could refer to different columns such as ”donator_name” or ”School_name,” the system should ask for clarification (e.g., ”What did you mean by…?”) or inquire about which columns to return, while also generating SQL queries based on these interpretations.
  2. 2.
    Answerable: These questions can be directly answered using SQL queries based on the available database information. The system must generate an accurate SQL query reflecting the user’s request and execute it to retrieve the required data. For instance, if the user asks for the ”id of Glenn” and it is clear that ”Glenn” refers to the ”donator_name” column, the system should generate the appropriate SQL query.
  3. 3.
    Unanswerable: These questions pertain to information that is not present in the database or exceeds the system’s operational scope, such as when they involve external knowledge or web search capabilities beyond the SQL system. The system should acknowledge its inability to answer the question and explain the reason to the user. For example, a question about the ”country” of donors, if not present in the database, should be identified by the system as unanswerable.
  4. 4.
    Improper: These questions are irrelevant to the database or pertain to everyday conversation that does not require an SQL response. The system should recognize these questions and respond appropriately, without attempting to generate an SQL query. For example, a casual ”thank you” from the user should be met with a simple ”you’re welcome” rather than an SQL statement.
In this task, the model is required to perform two key functions: first, to accurately determine the type of each question; and second, to attempt resolution and provide answers for questions categorized as ambiguous and answerable. This dual focus ensures that the model classifies questions effectively and actively generates appropriate SQL queries for complex scenarios.
3.2 Construction
The MMSQL data originates from two sources: samples refined from CoSQL and new samples generated by QDA-SQL [15]. QDA-SQL employs CoT to guide the generation of new samples, helping the LLM produce multi-turn Q&A pairs by reasoning step-by-step and constructing intermediate thoughts leading to the final answer. This uses iterative processes, random combinations of context relationships, and question types, followed by a refinement process to guide Gemini Pro in generating diverse multi-turn, multi-type datasets, ensuring each sample adheres to our defined question types. The samples generated by QDA-SQL demonstrated higher natural language annotation quality, with a 62% win rate, and included more complex text-to-SQL examples compared to the original dataset. This resulted in better performance with Enhanced Data compared to the original dataset when used to train CodeLlama. We screened these initial samples to ensure they met high standards of question difficulty, classification accuracy, and diversity, resulting in a high-quality training set. Necessary modifications and optimizations were made to the original datasets, transforming some instances into a curated MMSQL test set. Unanswerable and ambiguous questions are sometimes not easily distinguishable, so we ensured these characteristics were prominent enough during the manual annotation of the test set.
3.3 Dataset Analysis
Table 1 summarizes the statistics of MMSQL, compared to two other multi-turn or multi-type datasets. A distinctive feature of MMSQL is that it includes four types of questions. MMSQL offers a significantly larger number of dialogues and features a higher average number of turns per dialogue, providing abundant multi-turn training data that covers a wide range of query types.
Table 1: Comparison of multi-turn or multi-type text-to-SQL datasets.
@@ 3.3 Dataset Analysis
3.4 Evaluation Metrics
@@
To evaluate the quality of a model’s responses, we use an LLM-assisted rating method assessing Utility, Accuracy, Completeness, Clarity, and Relevance on a 0-to-2 scale, with a maximum score of 10. This approach surpasses traditional metrics like BLEU, ROUGE, and BERT-Score [32, 33]. We employ GPT-4o-mini due to its strong alignment with human judgment [34, 35, 36] and enhance evaluation calibration using the Multiple Evidence Calibration methodology [37]. To ensure alignment with human judgment, three experts scored each sample, and their scores were averaged. Pearson, Spearman, and Kendall’s Tau correlations were calculated, confirming a strong positive correlation between LLM and human scores, as shown in Table 2, demonstrating its reliability.
Table 2: Correlation analysis between human and GPT-4 ratings for responses to different types of questions where SQL answer is not applicable.
@@
Una. 0.86 5.44e-07 0.57 7.63e-03 0.54 8.76e-03
Amb. 0.96 4.07e-08 0.88 3.09e-05 0.81 3.86e-04
Imp. 0.71 3.46e-23 0.62 1.07e-16 0.61 7.14e-14
[…]
Table 3: Performance of the models on the MMSQL test set in complex multi-turn Text-to-SQL tasks.
Model TDEX IEX EX EM RQS Ans. Una. Amb. Imp. F1
Prec Rec Prec Rec Prec Rec Prec Rec
GPT-4 Turbo 67.0 30.2 70.0 51.0 5.80 90.3 89.8 25.9 70.0 56.9 38.4 100.0 98.0 68.2
GPT-3.5 Turbo 64.1 25.5 69.6 47.3 4.74 88.8 89.2 16.0 65.0 64.3 20.9 100.0 96.0 61.1
Gemini-1.5 Flash 65.8 30.1 70.0 52.3 4.03 85.9 95.8 26.3 50.0 58.3 8.1 100.0 95.4 59.3
Llama3-70B 62.8 22.8 66.4 47.4 3.86 84.9 95.2 27.5 55.0 80.0 9.3 100.0 92.7 59.8
Llama3-8B 64.0 20.1 66.1 45.7 4.55 88.3 93.2 21.4 60.0 83.3 23.3 100.0 96.7 64.2
SQLCoder-8B 30.7 24.8 63.2 31.0 3.43 84.6 99.6 77.8 35.0 0.0 0.0 100.0 98.0 59.7
Codellama-7B 30.7 3.4 27.2 21.7 5.09 93.9 16.5 4.3 85.0 96.6 66.3 56.8 100.0 46.9
Mistral-7B-v0.2 26.4 0.7 13.7 11.2 4.37 82.1 57.6 4.7 55.0 100.0 50.0 79.1 77.5 55.3
[…]
00footnotetext: The best score in each column is highlighted in bold, and the second-highest score is underlined. Additionally, the models’ abilities to recognize Answerable, Unanswerable, Ambiguous, and Improper questions are evaluated, with precision and recall provided for each category. F1 represents the average F1 score across all categories. All in % except for average RQS.
3.5 Baselines
We evaluate eight popular LLMs’ performance, including close-source and open-source options. The close-source LLMs evaluated are GPT-4 Turbo [^1] https://platform.openai.com/docs/models/GPT-4-Turbo-and-gpt-4, GPT-3.5 Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, and Gemini-1.5 Flash [^3] https://deepmind.google/technologies/gemini. The open-source LLMs include Llama-3 (70B and 8B) [^4] https://llama.meta.com/llama3, Llama-3-SQLCoder-8B [^5] https://defog.ai (a Llama-3 model specifically trained in the text-to-SQL domain), Codellama-7B [^6] https://www.llama.com/docs/integration-guides/meta-code-llama, and Mistral-7B-v0.2777https://mistral.ai. See Appendix A for experimental details.
3.6 Result of MMSQL Evaluation
Table 3 presents the performance of the evaluated LLMs in the zero-shot evaluation on multi-turn text-to-SQL tasks of the MMSQL test set. This test set encompasses four types of questions. We assessed the models based on overall performance (TDEX), SQL generation performance (EX and EM), and their ability to recognize different types of questions. We provide a detailed discussion of the results in the following sections.

### v2

# Harbin, China 150040

@@ Abstract
Recent advancements in large language models (LLMs) have significantly improved text-to-SQL systems. However, most datasets and LLM-based methods tend to focus narrowly on SQL generation, often neglecting the complexities inherent in real-world conversational queries. This oversight can result in unreliable responses, particularly for ambiguous questions that cannot be directly addressed with SQL. To address this gap, we propose MMSQL, a comprehensive test suite designed to evaluate LLMs’ question classification and SQL generation capabilities by simulating real-world scenarios with diverse question types and multi-turn Q&A interactions. Utilizing MMSQL, we assessed the performance of popular LLMs, including both open-source and closed-source models, and identified key factors influencing their performance in these contexts. Furthermore, we introduce an LLM-based multi-agent framework that employs specialized agents to identify question types and determine appropriate answering strategies. Experimental results demonstrate that this method effectively enhances baseline models’ ability to handle diverse question types in conversational scenarios. Our approach simultaneously considers multiple question types and multi-turn interactions, providing a new, realistic perspective, and offering a valuable advancement toward more reliable and versatile text-to-SQL systems.
Index Terms:
@@
I Introduction
Text-to-SQL bridges natural language and SQL queries, empowering non-technical users to access and analyze data without mastering complex SQL knowledge. The advent of LLMs, with their remarkable capacity for following instructions, has transformed the text-to-SQL domain. LLM-based methods have delivered remarkable outcomes across various text-to-SQL tasks [1, 2], while the assessment of their robustness is increasingly drawing attention. Current research often assumes that user queries are unambiguous, focusing on expanding the scope of databases and increasing the complexity of questions to elevate the difficulty of question-to-SQL mapping, thereby pushing the boundaries of text-to-SQL systems. [3, 4, 5, 6]. However, in practical scenarios, user queries are inherently dynamic and uncertain [7]. Assuming that all user queries can be answered may lead to the generation of hallucinated content and unreliable predictions, thereby undermining the dependability of these systems [8, 9].
Observing real-world conversations, it is common for users to engage in multi-turn dialogues with interrelated questions that do not readily translate into SQL. In such scenarios, relying on single-turn or assumptive responses can lead to suboptimal performance, as it may not fully address the user’s needs or context [7, 10, 11]. Figure 1 illustrates examples of responses to three common types of such questions: (1) Unanswerable: If the database lacks data, such as the flight departure times from an airport queried in turn 1, the system needs to explain its inability to provide an answer. (2) Ambiguous: In turn 3, the query about the flight number of Delta Airlines might refer to conditions from turn 1 and turn 2: "from airport APG", or it might simply mean all flight numbers operated by Delta Airlines. The system needs to recognize this ambiguity and respond accordingly to avoid providing incorrect SQL responses. (3) Improper: Questions unrelated to the database content, such as casual conversation in turn 4, should not generate SQL-based responses. These examples highlight the necessity for reliable models capable of navigating various types of complex conversational dynamics. There is a need for more robust models that can effectively manage such challenges.
@@ I Introduction
Existing studies have only sporadically addressed the effective assessment and handling of multi-type questions. Most text-to-SQL research focuses on achieving high accuracy for single-type or single-round user questions [4, 6, 12, 13], often overlooking the need to develop systems capable of multi-turn dialogues that handle a variety of question types [3]. This limited scope can result in incorrect responses, especially for those questions that do not necessitate SQL solutions. It has been observed that even the most sophisticated LLMs struggle to effectively address these challenges [9]. Although recent studies have acknowledged the challenges of ambiguous and unanswerable questions [7]. However, existing solutions often resort to simple abstention when faced with vague or conversational elements, lacking granular designs to manage various question types, leading to passive model responses and hindering practical applicability and robustness [10, 14, 15]. In summary, while individual studies have made some progress in evaluating and enhancing the reliability of text-to-SQL systems, these efforts fall short of a comprehensive solution to the problem.
In response to these challenges, we introduce the Multi-type and Multi-turn text-to-SQL test suite (MMSQL), a comprehensive benchmark engineered to evaluate the proficiency of LLMs in handling multi-turn text-to-SQL tasks across diverse question types. Utilizing MMSQL, we have undertaken an exhaustive performance analysis of leading LLMs, encompassing open-source and closed-source models. This analysis provides detailed insights, highlighting the promise of open-source models and reveals the suboptimal performance of models on specific question types and the varying effectiveness of strategies designed to resolve ambiguity. As an initial attempt to address these challenges, we also propose an innovative LLM-based multi-agent framework, inspired by top-performing models and informed by these insights. This framework is anchored by a core Question Detector and Question Decomposer, tasked with identifying question types and determining appropriate answering strategies. It could provide potential SQL queries for ambiguous questions, accounting for multiple interpretations. Furthermore, the framework includes two supportive agents: the Schema Selector, which identifies and provides the essential subset of a database schema, and the SQL Refiner, which is dedicated to refining SQL queries. Our experiments, conducted on the MMSQL benchmark, have shown that this framework yields a significant performance improvement.
@@
II Related Work
II-A Text-to-SQL
@@
Text-to-SQL involves transforming natural language queries into SQL, representing a crucial area of natural language processing with many practical applications. Essential datasets include single-turn datasets such as WikiSQL [16], Spider[17], and BIRD [3], as well as multi-turn datasets such as CHASE [13], SParC [12], and CoSQL [18]. These datasets serve as benchmarks for evaluating the performance of text-to-SQL systems. In recent years, the emergence of pre-trained language models (PLMs) has driven significant progress in this domain. Models like SQLova [19], PICARD [20], and RAT-SQL [21] have achieved impressive results. However, despite their effectiveness, these specialized models require considerable refsources and time for training. LLMs have revolutionized numerous NLP tasks, including text-to-SQL, surpassing previous models and establishing a new paradigm [22]. Recent studies have focused on enhancing prompt design and developing complex, multi-stage frameworks that leverage LLMs to enhance performance [23]. Notably, LLM-based frameworks like DIN-SQL [24], which utilize chain-of-thought reasoning [25], decompose the problem into simpler sub-problems. MAC-SQL [6] and CHASE-SQL [5] employ multi-agent collaborative frameworks, demonstrating significant advancements in this field.
@@
Recent studies highlight the inherent diversity and ambiguity of natural language, noting that not all user queries can be effectively translated into accurate SQL as the answer in practical applications [7, 10, 26]. Datasets such as TriageSQL [14], NoisySP [7], and CoSQL have incorporated question type detection as part of the models’ tasks. Previous work, TrustSQL [10] demands that models opt for abstention when faced with diverse unanswerable questions. AMBROSIA [9] and [27] propose testing text-to-SQL systems by parsing ambiguous questions into multiple potential SQL queries, thereby enhancing model usability in real-world scenarios. [27] improves beam search on T5 [28] to provide possible SQL responses for ambiguous questions and evaluate their effectiveness. Additionally, [15] employs questions enhanced dialogue augmentation to train open-source LLMs to identify question types. These studies focus solely on the model’s ability to recognize question types or address specific problems, which may result in unhelpful abstentions or hallucinations due to specialized question-and-answer tailoring [29]. Moreover, most work is limited to single-turn settings, whereas real interactions often involve ambiguity that must be addressed in multi-turn contexts. These limitations motivate our work, which evaluates and enhances LLMs for multi-turn dialogues with multiple question types.
II-B LLM-based Agents
@@
LLM-based agents have long captured the attention of researchers in both academia and industry [30]. With the expansion of web knowledge, LLMs are increasingly demonstrating intelligence levels comparable to humans. This evolution has sparked a growing interest in creating autonomous agents driven by LLMs. AutoGPT [31] and MetaGPT [32] enhance AI models by integrating a range of useful tools, allowing developers to create adaptable, conversational agents that can function in various modes by leveraging LLMs, human input, and other tools to complete tasks. In the realm of text-to-SQL parsing, frameworks such as MAC-SQL [6], MAG-SQL [33], and CHASE-SQL [5] employ multiple LLM-based agents to interpret SQL queries collaboratively. This can typically be summarized as a three-step process: retrieving relevant values, generating SQL, and refining the SQL output. These multi-agent systems have achieved leading results on the BIRD benchmark [3]. However, existing methodologies have yet to incorporate pipeline designs specifically tailored to manage multiple types of questions concurrently within interactive environments. Drawing inspiration from these advancements, we developed a multi-agent framework incorporating a Question Detector Agent, designed to effectively manage the complexity and variety of user queries encountered in real-world applications.
III MMSQL: Multi-type and Multi-turn Text-to-SQL Test Suite
TABLE I: Comparison of multi-turn or multi-type text-to-SQL datasets.
@@ III MMSQL: Multi-type and Multi-turn Text-to-SQL Test Suite
III-A Task Formulation
  1. 1.
    Unanswerable: These questions pertain to information that is not present in the database or exceeds the system’s operational scope, such as when they involve external knowledge or web search capabilities beyond the SQL system. The system should acknowledge its inability to answer the question and explain the reason to the user. For example, a question about the airport’s flight schedule each day, which is not present in the database, should be identified by the system as unanswerable.
  2. 2.
    Answerable: These questions can be directly answered using SQL queries based on the available database information. The system must generate an accurate SQL query reflecting the user’s request and execute it to retrieve the required data. For example, in Turn 2, when asked, "How many airlines have flights departing from here?" the system correctly generates the SQL query by interpreting "here" within the given context to find airlines with flights departing from the specified location.
  3. 3.
    Ambiguous: These questions include terms that might correspond to multiple columns or values in the database, causing potential ambiguity in SQL query generation. In a multi-turn context, the system should identify and address this ambiguity by seeking clarification from the user. For example, in Turn 3, when asked, "What is the flight number of Delta Airlines?" the system should recognize the potential ambiguity—whether the user is asking for all flights or specific ones from a location like APG. It should ask for clarification (e.g., "Are you asking about all flights or specific ones?") and simultaneously propose possible SQL queries based on these interpretations. This approach ensures the system can adapt and refine its responses through ongoing dialogue.
  4. 4.
    Improper: These questions are irrelevant to the database or pertain to everyday conversation that does not require an SQL response. The system should recognize these questions and respond appropriately, without attempting to generate an SQL query. For example, a casual "thank you" from the user should be met with a simple "You’re welcome!" rather than an SQL statement.
In this task, the model is required to perform two key functions: first, to accurately determine the type of each question; and second, to attempt resolution and provide answers for questions categorized as ambiguous and answerable. This dual focus ensures that the model classifies questions effectively and actively generates appropriate SQL queries for complex scenarios. For questions deemed improper or unanswerable, the model aims to provide the most helpful natural language responses possible.
III-B Construction
The MMSQL data comes from two sources: refined samples from CoSQL and new samples generated by QDA-SQL (Questions Enhanced Dialogue Augmentation for Multi-turn Text-to-SQL) [15]. QDA-SQL uses Chain of Thought (CoT) [25] to guide new sample generation through step-by-step reasoning. This involves multi-turn Q&A generation, and LLM-based refinement to help Gemini-Pro generate diverse samples. Each sample aligns with predefined question types. Appendix B provides further details on its effectiveness and design.
Figure 2: Distribution of question types in the curated MMSQL dataset, highlighting the proportions of Answerable, Ambiguous, Improper, and Unanswerable questions in both the training set and the test set.
We carefully screened initial samples to ensure high standards of difficulty, classification accuracy, and diversity, resulting in a refined dataset. Modifications transformed some instances into a curated MMSQL test set, with a focus on clearly distinguishing unanswerable and ambiguous questions during manual annotation. The dataset includes 6,493 training rounds and 149 test rounds of dialogue. Figure 2 shows the distribution of question types, highlighting that the training set covers all four types sufficiently, while the test set reflects a more realistic distribution [7]. Table I compares MMSQL with other multi-turn or multi-type datasets, MMSQL uniquely includes four question types and offers more dialogues with a higher average number of turns per dialogue, providing extensive multi-turn training data across diverse query types.
III-C Evaluation Metrics
@@
To evaluate the quality of a model’s responses, MMSQL utilizes an LLM-assisted rating method assessing Utility, Accuracy, Completeness, Clarity, and Relevance on a 0-to-2 scale, with a maximum score of 10. This approach, where LLMs directly score responses, has been demonstrated to be superior in assessing answer quality compared to traditional metrics like BLEU, ROUGE, and BERT-Score. [34, 35]. The detailed prompts used for this evaluation are provided in Appendix C. We employ GPT-4o-mini due to its strong alignment with human judgment [36, 37] and enhance evaluation calibration using the Multiple Evidence Calibration methodology [38]. To ensure alignment with human judgment, three experts scored each of 100 samples, and their scores were averaged. Pearson, Spearman, and Kendall’s Tau correlations were calculated, confirming a strong positive correlation between LLM and human scores, as shown in Table II, demonstrating its reliability.
TABLE II: Correlation analysis between human and GPT-4 ratings for responses to different types of questions where SQL answer is not applicable.
@@
Una. 0.86 5.4e-07 0.57 7.6e-03 0.54 8.7e-03
Amb. 0.96 4.0e-08 0.88 3.0e-05 0.81 3.8e-04
Imp. 0.71 3.4e-23 0.62 1.0e-16 0.61 7.1e-14
III-D Baselines
We evaluate eight popular LLMs’ performance, including close-source and open-source options. The close-source LLMs evaluated are GPT-4 Turbo [^1] https://platform.openai.com/docs/models/GPT-4Turbo-and-gpt-4, GPT-3.5 Turbo [^2] https://platform.openai.com/docs/models/gpt-3-5-turbo, and Gemini-1.5 Flash [^3] https://deepmind.google/technologies/gemini. The open-source LLMs include Llama-3 (70B and 8B) [^4] https://llama.meta.com/llama3, Llama-3-SQLCoder-8B [^5] https://defog.ai (a Llama-3 model specifically trained in the text-to-SQL domain), Codellama-7B [^6] https://www.llama.com/docs/integration-guides/meta-code-llama, and Mistral-7B-v0.2777https://mistral.ai.
III-E Result of MMSQL Evaluation
Table III and Table V showcases the performance of the evaluated LLMs in a zero-shot setting on the multi-turn text-to-SQL task, covering four types of questions from the MMSQL test set. We assessed the models based on overall performance (TDEX), SQL generation performance (EX and EM), and their capability to recognize different question types. A detailed discussion of the results is provided in the following sections.
TABLE III: Performance metrics of the models on the MMSQL test set, with all values expressed in percentages, except for the RQS.
Model TDEX IEX EX EM RQS
GPT-4 Turbo 67.0 30.2 70.0 51.0 5.80
GPT-3.5 Turbo 64.1 25.5 69.6 47.3 4.74
Gemini-1.5 Flash 65.8 30.1 70.0 52.3 4.03
Llama3-70B 62.8 22.8 66.4 47.4 3.86
Llama3-8B 64.0 20.1 66.1 45.7 4.55
SQLCoder-8B 30.7 24.8 63.2 31.0 3.43
Codellama-7B 30.7 3.4 27.2 21.7 5.09
Mistral-7B-v0.2 26.4 0.7 13.7 11.2 4.37